40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12099     Entry time: Fri Apr 29 00:55:46 2016     In reply to: 12095     Reply to this: 12104
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: green PDH locked to Xarm 

Using the modulation frequency suggested here, I hooked up the PDH setup at the X-end and succeeded in locking the green to the X arm. I then rotated the HWP after the green Faraday to maximize TRX output, which after a cursory alignment optimization is ~0.2 (I believe we were used to seeing ~0.3 before the end laser went wonky). Obviously much optimization/characterization remains to be done. But for tonight, I am closing the PSL and EX laser shutters and applying first contact to the window once more courtesy more PEEK from Koji's lab in W Bridge. Once this is taken care of, I can install the Oplev tomorrow, and then set about optimizing various things in a systematic way.. MC autolocker has also been disabled...

Side note: for the IR Transmon QPD, we'd like a post that is ~0.75" taller given the difference in beam height from the arm cavity and on the endtable. I will put together a drawing for Steve tomorrow..

ELOG V3.1.3-